Question 30
ProbabilityIf two fair coins are tossed, what is the probability of getting at least one head?
- A.1/4
- B.1/2
- C.1
- D.2/3
- E.3/4
Show the answer and explanation
Answer: E (3/4)
Explanation
The only outcome with no head is TT, with probability ¼. So P(at least one head) = 1 − ¼ = ¾.